OpenStudy (anonymous):

A rectangle has perimeter 40 cm. Let the length of one side be x. Then write a function A(x) that represents that area of the rectangle

OpenStudy (anonymous):

if you call the other side \(w\) then \(2w+2x=40\) and \[A=wx\] solve the first equation for \(w\) and get \[2w+2x=40\\ 2w=40-2x\\ w=20-x\] then substitute back in to \[A=wx\] and get \[A(x)=(20-x)x\] or if you prefer \[A(x)=20x-x^2\]

OpenStudy (anonymous):

Do you know how I would graph this on a graphing calculator by chance?